# Neural Flow Diffusion Models: Learnable Forward Process for Improved Diffusion Modelling

### v1

@@ Abstract
Conventional diffusion models typically relies on a fixed forward process, which implicitly defines complex marginal distributions over latent variables. This can often complicate the reverse process’ task in learning generative trajectories, and results in costly inference for diffusion models. To address these limitations, we introduce Neural Flow Diffusion Models (NFDM), a novel framework that enhances diffusion models by supporting a broader range of forward processes beyond the fixed linear Gaussian. We also propose a novel parameterization technique for learning the forward process. Our framework provides an end-to-end, simulation-free optimization objective, effectively minimizing a variational upper bound on the negative log-likelihood. Experimental results demonstrate NFDM’s strong performance, evidenced by state-of-the-art likelihood estimation. Furthermore, we investigate NFDM’s capacity for learning generative dynamics with specific characteristics, such as deterministic straight lines trajectories. This exploration underscores NFDM’s versatility and its potential for a wide range of applications.
diffusion, generative models, variational inference
@@ 1 Introduction
Generative models are a versatile class of probabilistic machine learning models, which finds applications across diverse fields such as art, music, medicine, and physics (Tomczak, 2022; Creswell et al., 2018; Papamakarios et al., 2021; Yang et al., 2022). Generative models excel in constructing probability distributions that accurately represent datasets, enabling them to produce new samples akin to the original data. Such capabilities make them ideal for tasks like enhancing training data through the generation of synthetic datasets (data augmentation) and uncovering hidden patterns and structures in an unsupervised learning setting.
@@
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of generative models are constructed by two key processes: the forward process and the reverse process. The forward process gradually corrupts the data distribution, transforming it from its original form to a noised state. The reverse process learns to invert corruptions of the forward process and restore the data distribution. This way, it learns to generate data from pure noise. Diffusion models have demonstrated remarkable results in various domains (Ho et al., 2022; Saharia et al., 2022; Tan et al., 2024; Watson et al., 2023; Trippe et al., 2023). Nevertheless, most existing diffusion models fix the forward process to be predefined Gaussian, which makes it unable to adapt to the task at hand or simplify the target for the reverse process. At the same time there are many works that demonstrate how modifications of the forward process improve performance in terms of generation quality (Nichol & Dhariwal, 2021; Vahdat et al., 2021; Daras et al., 2022), likelihood estimation (Kingma et al., 2021; Nielsen et al., 2023; Bartosh et al., 2023) or sampling speed (Lee et al., 2023; Pooladian et al., 2023; Tong et al., 2023).
@@
In this paper, we present Neural Flow Diffusion Models (NFDM), a framework that allows for the pre-specification and learning of latent variable distributions defined by the forward process. Unlike conventional diffusion models (Ho et al., 2020), which rely on a conditional Gaussian forward process, NFDM may accommodate any continuous (and learnable) distribution that can be expressed as an invertible mapping applied to noise. We also derive, and leverage, an end-to-end simulation-free optimization procedure, that minimizes a variational upper bound on the negative log-likelihood (NLL).
@@
We also propose an efficient neural network-based parameterization for the forward process, enabling it to adapt to the reverse process during training and simplify the learning of the data distribution. To demonstrate NFDM’s capabilities with a learnable forward process we provide experimental results on CIFAR-10, ImageNet 32 and 64, attaining state-of-the-art NLL results, which is crucial for many applications such as data compression (Ho et al., 2021; Yang & Mandt, 2022), anomaly detection (Chen et al., 2018b; Dias et al., 2020) and out-of-distribution detection (Serrà et al., 2019; Xiao et al., 2020).
@@
Leveraging the flexibility of NFDM, we further explore training with constraints on the reverse process to learn generative dynamics with specific properties. As a case study, we discuss a curvature penalty on the deterministic generative trajectories. Our empirical results indicate improved computational efficiency compared to baselines on synthetic datasets as well as MNIST, CIFAR-10, and downsampled ImageNet.
@@
    We develope an end-to-end optimization procedure that minimizes an upper bound on the negative log-likelihood in a simulation-free manner.
@@
    We show how NFDM can be used in learning generative processes with specific properties, exemplified by dynamics with straight line trajectories, where NFDM leads to significantly faster sampling speeds and enhanced generation quality with fewer sampling steps.
@@ 2 Background
Due to the linearity of $\tilde{f}^{F}$, we can reconstruct the conditional marginal distribution $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}I)$. Typically, the conditional distributions evolve from some low variance distribution $q(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{z}_{0}-\mathbf{x})$ to a unit Gaussian $q(\mathbf{z}_{1}|\mathbf{x})\approx\mathcal{N}(\mathbf{z}_{1};0,I)$. This forward process is then reversed by starting from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and following the reverse SDE (Anderson, 1982):
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}(\mathbf{z}_{t},t)dt+g(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (2)
$\tilde{f}^{B}(\mathbf{z}_{t},t)$ $=\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}).$ (3)
Here, $\bar{\mathbf{w}}$ denotes a standard Wiener process where time flows backwards. Diffusion models approximate this reverse process by learning $\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t})$, known as the score function, through a $\lambda_{t}$-weighted score matching loss:
$\underset{u(t)q(\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t})\big{\|}_{2}^{2}\right],$ (4)
where $u(t)$ represents a uniform distribution over the interval $[0,1]$, and $s_{\theta}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ is a learnable approximation. This loss can be reformulated as a more tractable denoising score matching loss Vincent (2011):
$\underset{u(t)q(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\big{\|}_{2}^{2}\right].$ (5)
With a learned score function $s_{\theta}(\mathbf{z}_{t},t)$, one can generate a sample from the reverse process by first sampling from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and then simulating the reverse SDE:
$d\mathbf{z}_{t}$ $=\big{[}\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)s_{\theta}(\mathbf{z}_{t},t)\big{]}dt+g(t)d\bar{\mathbf{w}},$ (6)
resulting a sample $\mathbf{z}_{0}\sim p_{\theta}(\mathbf{z}_{0})\approx q(\mathbf{z}_{0})\approx q(\mathbf{x})$.
Diffusion models possess several important properties. For example, for specific $\lambda_{t}$, the objective (eq. 5) can be reformulated (Song et al., 2021) as an Evidence Lower Bound (ELBO) on the model’s likelihood. Furthermore, the minimization of denoising score matching (eq. 5) is a simulation-free procedure. This means that simulating either the forward or reverse processes through its SDE is not necessary for sampling $\mathbf{z}_{t}$, nor is it necessary for estimating the gradient of the loss function. Instead, we can directly sample $\mathbf{z}_{t}\sim q(\mathbf{z}_{t}|\mathbf{x})$. The simulation-free nature of this approach is a crucial aspect for efficient optimization.
@@
This implies that we can sample from diffusion models deterministically, allowing the use of off-the-shelf numerical ODE solvers for sampling, which may improve the sampling speed compared to stochastic sampling that requires simulating an SDE. Additionally, deterministic sampling enables us to compute densities by treating the model as a continuous normalizing flow, as detailed in Chen et al. (2018a); Grathwohl et al. (2018).
@@ 3 Neural Flow Diffusion Models
Diffusion models can be viewed as a specific type of hierarchical Variational Autoencoders (VAEs) (Kingma & Welling, 2013; Rezende et al., 2014), where the latent variables are inferred through a pre-specified scaling of the data point and injection of Gaussian noise. However, this formulation limits diffusion models in terms of the flexibility of their latent space, and makes learning of the reverse process more challenging. To address this limitation, we propose a generalized form of data and noise transformations that enables the definition and learning of a broad range of distributions in the latent space beyond the conventional Gaussian.
In this section, we introduce Neural Flow Diffusion Models (NFDM)– a framework that generalizes conventional diffusion models. The key idea in NFDM is to define the forward process implicitly via a learnable transformation $F_{\varphi}(\varepsilon,t,\mathbf{x})$. This lets the user define a broad range of continuous time and data dependent forward process. Importantly, NFDM retains crucial properties of conventional diffusion models, like likelihood-based and simulation-free training. Previous diffusion models emerge as special cases when the data transformation is linear, time-independent, and/or additive Gaussian. The connections between NFDM and existing models are discussed in Section 6.
@@ 3.1 Forward process
The forward process defines a target for the reverse process. It describes the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ in the latent space that the reverse process tries to match.
We approach the forward process constructively. First, we define a conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for $t\in[0,1]$. Then, we introduce the corresponding conditional ODE that together with the initial distribution $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ matches the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. Finally, we define a conditional SDE that implicitly defines distribution over trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, with the same marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
Forward Marginal Distribution. We characterize the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ of the forward process trough a function of injected noise $\varepsilon$, time $t$, and the data $\mathbf{x}$:
@@
where $F_{\varphi}:\mathbb{R}^{D}\times[0,1]\times\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$ and $\varepsilon\sim q(\varepsilon)=\mathcal{N}(\varepsilon;0,I)$. This function transforms a sample from a distribution $q(\varepsilon)$ into $\mathbf{z}_{t}$, conditional on the data point $\mathbf{x}$ and time step $t$, implicitly defining the conditional distribution of latent variables $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. Additionally, eq. 9 facilitates direct and efficient sampling from $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ through $F_{\varphi}$.
Conditional ODE. Assuming that $F_{\varphi}$ is differentiable with respect to $\varepsilon$ and $t$ and invertible with respect to $\varepsilon$, fixing specific values of $\mathbf{x}$ and $\varepsilon$ and varying $t$ from 00 to $1$ results in a smooth trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}_{1}$. Differentiating these trajectories over time yields a velocity field corresponding to the conditional distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, thereby defining a conditional ODE:
$d\mathbf{z}_{t}$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt,\quad\textrm{where}$ (10)
Therefore, if we sample $\mathbf{z}_{0}\sim q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ and solve the ODE in eq. 10 until time $t$, we have a sample $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
The time derivative of $F_{\varphi}$ may be calculated using automatic differentiation tools like PyTorch (Paszke et al., 2017) or JAX (Bradbury et al., 2018). Specifically, we can compute a Jacobian Vector Product (JVP) with a unit one-dimensional vector. Unlike standard backpropagation, which corresponds to a Vector Jacobian Product (VJP), JVP does not necessitate $\mathcal{O}(D)$ backpropagation passes, enhancing scalability.
Conditional SDE. The function $F_{\varphi}$ and the distribution of the noise $q(\varepsilon)$ together defines $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$. To completely define the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, we introduce a conditional SDE that starts from sample $\mathbf{z}_{0}$ and runs forward in time.
With access to both the ODE and score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, the SDE (Song et al., 2020b) corresponding to the distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, is:
@@
Here, $g_{\varphi}:[0,1]\mapsto\mathbb{R}_{+}$ is a scalar function, and $\mathbf{w}$ represents a standard Wiener process. It is important to note, that $g_{\varphi}$ only influences the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, not the marginal distributions. The Conditional SDE in eq. 12 has the same conditional marginal distribution, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, for any $g_{\varphi}$.
The score function of $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is:
$\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\nabla_{\mathbf{z}_{t}}\left[\log q(\varepsilon)+\log\left|J_{F}^{-1}\right|\right]$ (13)
$\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})\quad\textrm{and}\quad J_{F}^{-1}=\frac{\partial F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})}{\partial\mathbf{z}_{t}}.$ (14)
While the log-determinant of the Jacobian matrix $J_{F}^{-1}$ is not generally available in explicit form, it can be analytically evaluated for certain transformations or specific architectures of the function $F_{\varphi}$, like RealNVP style architectures (Dinh et al., 2016; Kingma & Dhariwal, 2018). The parameterization of $F_{\varphi}$ is further discussed in Section 3.5.
Given access to the log probability $\log q(\varepsilon)$ and the log-determinant of $J_{F}^{-1}$, automatic differentiation tools enable the calculation of gradients and the score function. Unlike the time derivative, regular backpropagation can be utilized here, ensuring computationally efficient score function evaluation.
3.2 Reverse process
To define the reverse (generative) process, we specify a reverse SDE that starts from $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$ and runs backwards in time. To do so we first introduce a conditional reverse SDE that reverses the conditional forward SDE (eq. 12). Following (Anderson, 1982), we define:
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w},\quad\textrm{where}$ (15)

### v2

@@ Abstract
Conventional diffusion models often rely on a fixed forward process, which implicitly defines complex marginal distributions over latent variables. This can often complicate the reverse process’ task in learning generative trajectories, and results in costly inference for diffusion models. To address these limitations, we introduce Neural Flow Diffusion Models (NFDM), a novel framework that enhances diffusion models by supporting a broader range of forward processes beyond the standard linear Gaussian. We also propose a novel parameterization technique for learning the forward process. Our framework provides an end-to-end, simulation-free optimization objective, effectively minimizing a variational upper bound on the negative log-likelihood. Experimental results demonstrate NFDM’s strong performance, evidenced by state-of-the-art likelihoods across a range of image generation tasks. Furthermore, we investigate NFDM’s capacity for learning generative dynamics with specific characteristics, such as deterministic straight lines trajectories, and demonstrate how the framework can be adopted for learning bridges between two distributions. The results underscores NFDM’s versatility and its potential for a wide range of applications.
@@ 1 Introduction
Diffusion models [55, 19] are a class of generative models constructed by two key processes: the forward process and the reverse process. The forward process gradually corrupts the data distribution, transforming it from its original form to a noised state. The reverse process learns to invert corruptions of the forward process and restore the data distribution. This way, the model learns to generate data from pure noise. Diffusion models have demonstrated remarkable results in various domains [20, 47, 62, 68, 64]. Nevertheless, most existing diffusion models fix the forward process to be predefined, usually linear, Gaussian which makes it unable to adapt to the task at hand or simplify the target for the reverse process. At the same time there is a growing body of work that demonstrates how modifications of the forward process improve performance in terms of generation quality [38, 65, 9], likelihood estimation [27, 39, 4] or sampling speed [31, 43, 63].
@@
In this paper, we present Neural Flow Diffusion Models (NFDM), a framework that allows for the pre-specification and learning of complex latent variable distributions defined by the forward process. Unlike conventional diffusion models [19], which rely on a conditional Gaussian forward process, NFDM can accommodate any continuous (and learnable) distribution that can be expressed as an invertible mapping applied to noise. We also derive, and leverage, a new end-to-end simulation-free optimization procedure for NFDM, that minimizes a variational upper bound on the negative log-likelihood (NLL).
@@
Furthermore, we propose an efficient neural network-based parameterization for the forward process, enabling it to adapt to the reverse process during training and simplify the learning of the data distribution. To demonstrate NFDM’s capabilities with a learnable forward process we provide experimental results on CIFAR-10, ImageNet 32 and 64, attaining state-of-the-art NLL results, which is crucial for many applications such as data compression [21, 70], anomaly detection [7, 13] and out-of-distribution detection [50, 69].
@@
Then, leveraging the flexibility of NFDM, we demonstrate how this framework can be applied to learn bridges between two distributions using the AFHQ dataset. Finally, we explore training with constraints on the reverse process to learn generative dynamics with specific properties. As a case study, we discuss curvature and obstacle avoidance penalties on the deterministic generative trajectories. Our empirical results indicate improved computational efficiency compared to baselines on CIFAR-10, downsampled ImageNet, and synthetic data.
@@
    We develop an end-to-end optimization procedure that minimizes an upper bound on the negative log-likelihood in a simulation-free manner.
@@
    We show how NFDM can be used in learning bridges and generative processes with specific properties, such as dynamics with straight line trajectories, leading to significantly faster sampling speeds and enhanced generation quality with fewer sampling steps.
@@ 2 Background
Due to the linearity of $\tilde{f}^{F}$, we can reconstruct the conditional marginal distribution $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}I)$. Typically, the conditional distributions evolve from some low variance distribution $q(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{z}_{0}-\mathbf{x})$ to a unit Gaussian $q(\mathbf{z}_{1}|\mathbf{x})\approx\mathcal{N}(\mathbf{z}_{1};0,I)$. This forward process is then reversed by starting from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and following the reverse SDE [3]:
$d\mathbf{z}_{t}=\tilde{f}^{B}(\mathbf{z}_{t},t)dt+g(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\tilde{f}^{B}(\mathbf{z}_{t},t)=\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}).$ (2)
Here, $\bar{\mathbf{w}}$ denotes a standard Wiener process where time flows backwards. Diffusion models approximate this reverse process by learning $\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t})$, known as the score function, through a $\lambda_{t}$-weighted denoising score matching loss:
$\underset{u(t)q(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}|\mathbf{x})\big{\|}_{2}^{2}\right],$ (3)
where $u(t)$ represents a uniform distribution over the interval $[0,1]$, and $s_{\theta}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ is a learnable approximation. With a learned score function $s_{\theta}(\mathbf{z}_{t},t)$, one can generate a sample from the reverse process by first sampling from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and then simulating the reverse SDE, resulting in a sample $\mathbf{z}_{0}\sim p_{\theta}(\mathbf{z}_{0})\approx q(\mathbf{z}_{0})\approx q(\mathbf{x})$:
$d\mathbf{z}_{t}$ $=\big{[}\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)s_{\theta}(\mathbf{z}_{t},t)\big{]}dt+g(t)d\bar{\mathbf{w}}.$ (4)
Diffusion models possess several important properties. For example, for a specific $\lambda_{t}$, the objective (eq. 3) can be reformulated [58] as an Evidence Lower Bound (ELBO) on the model’s likelihood. Furthermore, the minimization of denoising score matching (eq. 3) is a simulation-free procedure. This means that simulating either the forward or reverse processes through its SDE is not necessary for sampling $\mathbf{z}_{t}$, nor is it necessary for estimating the gradient of the loss function. Instead, we can directly sample $\mathbf{z}_{t}\sim q(\mathbf{z}_{t}|\mathbf{x})$. The simulation-free nature of this approach is a crucial aspect for efficient optimization.
@@
This implies that we can sample from diffusion models deterministically, allowing the use of off-the-shelf numerical ODE solvers for sampling, which may improve the sampling speed compared to stochastic sampling that requires simulating an SDE. Additionally, deterministic sampling enables us to compute densities by treating the model as a continuous normalizing flow, as detailed in [6, 16].
@@ 3 Neural Flow Diffusion Models
In diffusion models the forward process defines stochastic conditional trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ and the reverse process tries to match the marginal distribution of trajectories. This construction can be viewed as a specific type of hierarchical Variational Autoencoders (VAEs) [29, 44]. However, in conventional diffusion models the latent variables are inferred through a pre-specified linear combination of the data point and Gaussian noise. This formulation limits diffusion models in terms of the flexibility of their latent space, and makes learning of the reverse process more challenging. To address this limitation, we propose a generalized form of the forward process that enables the definition and learning of a broad range of distributions in the latent space. From a practical perspective, a more flexible forward process can simplify the task of learning the reverse process. From a theoretical perspective, learning of the forward process is analogous to learning the variational distribution in a hierarchical VAE, which gives a tighter bound on model’s NLL (see an extended discussion in Section B.1).
In this section, we introduce Neural Flow Diffusion Models (NFDM)– a framework that generalizes conventional diffusion models. The key idea in NFDM is to define the forward process’ conditional SDE implicitly via a learnable transformation $F_{\varphi}(\varepsilon,t,\mathbf{x})$ that defines the marginal distributions. This lets the user define a broad range of continuous time- and data-dependent forward processes, that the reverse process will learn to invert. Importantly, NFDM retains crucial properties of conventional diffusion models, like likelihood-based and simulation-free training. Previous diffusion models emerge as special cases when the data transformation is linear, time-independent, and/or additive Gaussian.
@@ 3.1 Forward Process
We approach the forward process constructively. The ultimate goal is a learnable distribution over trajectories, $\{\mathbf{z}(t)\}_{t\in[0,1]}$ given $\mathbf{x}$, realized by a conditional SDE constructed in three steps.
First, we (implicitly) define the conditional marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for $t\in[0,1]$ using $F_{\varphi}(\varepsilon,t,\mathbf{x})$. Then, we introduce the corresponding conditional ODE that together with an initial distribution $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ matches the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. Finally, we define a conditional SDE that defines a distribution over trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, with marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for each $t$.
Forward Marginal Distribution. We characterize the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ of the forward process trough a function that transforms a noise sample $\varepsilon$ into $\mathbf{z}_{t}$, conditional on the time step $t$ and data point $\mathbf{x}$:
@@
where $F_{\varphi}:\mathbb{R}^{D}\times[0,1]\times\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$ and $\varepsilon\sim q(\varepsilon)=\mathcal{N}(\varepsilon;0,I)$. This defines the conditional distribution of latent variables $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. Additionally, eq. 7 facilitates direct and efficient sampling from $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ through $F_{\varphi}$ in eq. 7.
Conditional ODE. We assume that $F_{\varphi}$ is differentiable with respect to $\varepsilon$ and $t$, invertible with respect to $\varepsilon$. Further, we assume that fixing specific values of $\mathbf{x}$ and $\varepsilon$ and varying $t$ from 00 to $1$ results in a smooth trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}_{1}$. Differentiating these trajectories over time yields a velocity field corresponding to the conditional distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, thereby defining a conditional ODE:
$d\mathbf{z}_{t}=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt,\quad\textrm{where}\quad f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})=\left.\frac{\partial F_{\varphi}(\varepsilon,t,\mathbf{x})}{\partial t}\right|_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}.$ (8)
Therefore, if we sample $\mathbf{z}_{0}\sim q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ and solve eq. 8 until time $t$, we have $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
The time derivative of $F_{\varphi}$ may be calculated efficiently using automatic differentiation tools like PyTorch [41] or JAX [5] (see details in Section B.2).
Conditional SDE. The function $F_{\varphi}$ and the distribution of the noise $q(\varepsilon)$ together defines $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. To completely define the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, we introduce a conditional SDE that starts from sample $\mathbf{z}_{0}$ and runs forward in time.
With access to both the ODE and score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, an SDE [60] with marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, is:
@@
Here, $g_{\varphi}:[0,1]\mapsto\mathbb{R}_{+}$ is a scalar function, and $\mathbf{w}$ represents a standard Wiener process. Note that $g_{\varphi}$ only influences the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$. The marginal distributions are the same, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, for any choice of $g_{\varphi}$.
The SDE in eq. 9 requires access to the conditional score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which in the general case can be computationally costly. However, for $F_{\varphi}$ that allow efficient evaluation of the log-determinant of the Jacobian matrix, the score function can be calculated efficiently. Examples of such $F_{\varphi}$ are functions linear with respect to $\varepsilon$ or RealNVP style architectures [14, 28]. The calculation of this score function is further discussed in Section B.3.
3.2 Reverse (Generative) Process
To define the reverse (generative) process, we specify a reverse SDE that starts from $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$ and runs backwards in time. To do so we first introduce a conditional reverse SDE that reverses the conditional forward SDE (eq. 9). Following [3], we define:
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (10)